# Leveraging Importance Sampling to Detach Alignment Modules from Large Language Models

### v1

@@ 2.1 Preliminary
Importance sampling estimates properties of a target distribution using samples from a different distribution, which is useful when direct sampling is difficult or of large variance. The method involves reweighting the samples to account for the differences between distributions:
@@ 2.3 Sequence-level Training
For training, we maximize this lower bound by emphasizing the term $\mathbb{E}_{{\bm{x}}\sim{\mathcal{S}},{\bm{y}}\sim P_{\mathrm{M}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]$, aligning it with the likelihood objective, while minimizing $-\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{S}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]$ as a surrogate for the overall loss.
@@ 2. Align: Each candidate is assigned an importance weight, serving as an aligning indicator:
It is important to note that the term $Q_{\theta}(y_{l}|y_{<l},{\bm{x}})$ is represented as a $\mathrm{Softmax}$ function in language models. This allows us to derive a shortcut implementation for secondary sampling similar to Nucleus Sampling, detailed in Appendix C, enabling the process to utilize the standard $\mathrm{Softmax}$ and sampling procedure for effective token selection.
@@ 3 Experimental Setup
Preference Optim. Anthropic-HH 169K $\geq$ 1 Preference Optimization
Training settings. We start the Proposal Module with the original pre-trained model and first conduct a warm-up phase to address distributional shifts. For supervised learning, we use thousands of examples to fine-tune the Proposal Module, enabling it to effectively generate the end-of-sequence token and appropriately conclude conversations. In the case of preference optimization, we follow the approach from [28] tp perform SFT using chosen responses. Prior to training, we sample from the Proposal Module across the entire dataset for supervised learning to create the training set. In contrast, the preference optimization allows us to train directly on preference labels, where the chosen response serves as the target and the rejected one acts as the proposal.
@@ 4 Experimental Results
Table 4: Performance comparison of Llama3 and Qwen2.5 RAM against baselines on the Anthropic-HH. Evaluation conducted using the AlpacaEval 2 framework with regards to helpfulness and harmlessness. "Ali." refers to the Aligner, and "R.A." refers to our Residual Aligner.
@@ 6 Related Works
Residual Correction for LLMs. Residual Energy-Based Models (Residual EBMs) [26, 7] enhance text generation by modeling the energy landscape to improve output coherence and control. They build on Energy-Based Models (EBMs) [16, 20, 30] by integrating globally normalized EBMs with local language models, refining a base distribution through energy-based adjustments to capture missed dependencies. The Aligner [18, 24] fine-tunes an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, stacking this onto the upstream model for corrected alignment. While effective in decoupling alignment from LLMs during training, the reliance on complete upstream responses introduces significant latency for the first token during inference. Additionally, the Aligner’s use of a reference response poses risks with out-of-distribution inputs. Our method leverages importance sampling to derive the residual alignment module, directly modeling conditional probabilities and mitigating out-of-distribution (OOD) risks from intermediate variables. Additionally, we introduce a token-level decoding strategy to achieve minimal first word latency, enhancing response robustness in practical applications.
@@ Appendix C Implementation Details of the Token-level Decoding
This reformulation simplifies implementation by allowing a logit pre-processor to be applied before the Residual Aligner computes the $\mathrm{Softmax}$. This pre-processor retains only the tokens sampled from the Proposal Module, setting the logits for other tokens to $-\mathrm{Inf}$, which is similar to the implementation of Nucleus Sampling. This adjustment enables the process to proceed through the standard $\mathrm{Softmax}$ and sampling procedure, allowing for effective token selection.
To mitigate performance degradation during the training of small Residual Aligner, $Q_{\theta}$, we only conduct secondary sampling when the distribution difference between the Proposal Module, $P_{\mathrm{M}}$, and the $Q_{\theta}$ is not significant. Specifically, we assess the difference using KL divergence $D_{KL}(P_{\mathrm{M}}\|Q_{\theta})$). If the KL divergence exceeds 0.1, indicating degradation of the Residual Aligner, we sample directly from the $P_{\mathrm{M}}$; otherwise, we apply the $Q_{\theta}$ for secondary sampling. Based on the Proposing-Aligning-Reducing (PAR) Sampling, we term it PAR_KL Sampling.

### v3

@@ 2.1 Preliminary
Importance sampling estimates properties of a target distribution using samples from a different distribution, which is useful when direct sampling is difficult. The method involves reweighting the samples to account for the differences between distributions:
@@ 2.3 Sequence-level Training
For training, we maximize this lower bound by emphasizing the term $\mathbb{E}_{{\bm{x}}\sim{\mathcal{S}},{\bm{y}}\sim P_{\mathrm{M}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]$, aligning it with the likelihood objective, while minimizing $\mathbb{E}_{({\bm{x}},{\bm{y}})\sim{\mathcal{S}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]$ as a surrogate for the overall loss.
@@ 2. Align: Each candidate is assigned an importance weight, serving as an aligning indicator:
This reformulation simplifies implementation by allowing a logit pre-processor to be applied before the Residual Aligner computes the $\mathrm{Softmax}$. This pre-processor retains only the tokens sampled from the Proposal Module, setting the logits for other tokens to $\mathrm{Inf}$, which is similar to the implementation of Nucleus Sampling. This adjustment enables the process to proceed through the standard $\mathrm{Softmax}$ and sampling procedure, allowing for effective token selection.
To mitigate potential performance degradation during the training of smaller Residual Aligners with fewer than 7 billion parameters, we implement secondary sampling only when the distribution difference between the Proposal Module, $P_{\mathrm{M}}$, and Residual Aligner, $Q_{\theta}$ is minimal. Specifically, we evaluate this difference using KL divergence, $D_{KL}(P_{\mathrm{M}}|Q_{\theta})$. If the KL divergence exceeds 0.1, indicating a degradation of the Residual Aligner, we sample directly from $P_{\mathrm{M}}$. Conversely, if the divergence is below this threshold, we utilize $Q_{\theta}$ for secondary sampling.
2.5 Variance Reduction
Importance sampling can result in high variance when the proposal distribution $P_{\mathrm{M}}$ poorly approximates the target distribution $P_{\mathcal{S}}$, often due to mismatches in support and probability. We assume that $P_{\mathrm{M}}$ supports $P_{\mathcal{S}}$ by leveraging the biased subset $\mathcal{S}$} from the training set, as discussed in 2.1, focusing specifically on probability mismatches.
The RAM introduces a learnable residual aligner $Q_{\theta}$ to adjust the alignment of $P_{\mathrm{M}}$ with $P_{\mathcal{S}}$, thereby reducing mismatch. By normalizing with the partition function $Z_{\theta}$, RAM ensures that $P_{\theta}$ remains a valid probability distribution. This normalization modifies the importance weight $W$, making it dependent on $Q_{\theta}$, which corrects deviations of $P_{\mathrm{M}}$ from $P_{\mathcal{S}}$, effectively smoothing large weights and minimizing variance.
During inference, RAM samples candidate tokens from Top-P regions of $P_{\mathrm{M}}$, maintaining higher $P_{\mathrm{M}}$ values and yielding lower importance weights $W$. We also employ Proposing-Aligning-Reducing Sampling, a self-normalized importance sampling method, to further reduce variance, despite introducing some bias.
In summary, we propose strategies in both training and inference to reduce variance and enhance stability with biased estimators.
@@ 3 Experimental Setup
Preference Optim. Anthropic-$H$ 169K $\geq$ 1 Preference Optimization
Training settings. We start the Proposal Module with the original pre-trained model and first conduct a warm-up phase to learn newly introduced special tokens (OOD tokens) in conversation tasks, such as "<|start_header_id|>", "<|end_header_id|>", "<|eot_id|>", etc., particularly within the Llama and Qwen model families. For supervised learning, we use thousands of examples to fine-tune the Proposal Module, enabling it to effectively generate the end-of-sequence token and appropriately conclude conversations. In the case of preference optimization, we follow the approach from [29] to perform SFT using chosen responses. Prior to training, we sample from the Proposal Module across the entire dataset for supervised learning to create the training set. In contrast, the preference optimization allows us to train directly on preference labels, where the chosen response serves as the target and the rejected one acts as the proposal.
@@ 4 Experimental Results
Table 4: Performance comparison of Llama3 and Qwen2.5 RAM against baselines on the Anthropic-$H$. Evaluation conducted using the AlpacaEval 2 framework with regards to helpfulness and harmlessness. "Ali." refers to the Aligner, and "R.A." refers to our Residual Aligner.
@@
We also conduct supplementary experiments to compare our model with Controlled Decoding (CD) [24], emphasizing both output quality and length in Appendix F.1. Additionally, we assess first-token latency in comparison to Aligner in Appendix F.2.
@@ 6 Related Works
Residual Correction for LLMs. Residual Energy-Based Models (Residual EBMs) [27, 7] enhance text generation by modeling the energy landscape to improve output coherence and control. They build on Energy-Based Models (EBMs) [16, 20, 31] by integrating globally normalized EBMs with local language models, refining a base distribution through energy-based adjustments to capture missed dependencies. Controlled Decoding (CD) [24] also takes the form of Residual EBMs which solve a KL-regularized RL objective to learn a prefix scorer for the reward that is used to steer the generation from a partially decoded path. The Aligner [18, 25] fine-tunes an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, stacking this onto the upstream model for corrected alignment. While effective in decoupling alignment from LLMs during training, the reliance on complete upstream responses introduces significant latency for the first token during inference. Additionally, the Aligner’s use of a reference response poses risks with out-of-distribution inputs. Our method leverages importance sampling to derive the residual alignment module, directly modeling conditional probabilities along with strategies during training and inference to reduce variance and enhance stability with biased estimators. Additionally, we introduce a token-level decoding strategy to achieve minimal first word latency, enhancing usability in practical applications.
@@ 7 Conclusion
Acknowledgments and Disclosure of Funding
This research is supported by Artificial Intelligence-National Science and Technology Major Project 2023ZD0121200 and the National Science Fund for Excellent Young Scholars under Grant 62222212.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state them in the introduction and abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have limitation section Appendix A
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper includes the complete set of assumptions for Proposition 2.1 along with a thorough and correct proof in Appendix B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the details in Appendix D
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide them in the supplemental materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the details in Section 3 and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We present the error bars for the experiments conducted in the Ablation Study in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See Section 3. We used a node with 8 80GB A800 Nvidia GPUs. We report details on batch sizes, number of gradient accumulation steps, and number of batches in Appendix D.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authores have reviewed the NeurIPS Code of Ethics and did not identify violations.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We describe it in the introduction and limitation sections on Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: N/A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all previous contributions (e.g., Hugging Face H4 for UltraChat, Anthropic for Anthropic-HH, OpenAI for TL;DR, Meta for LLama3, Alibaba for Qwen2.5).
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide an anonymous link/zip to our code which can be used for generating data and training models.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: N/A.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: N/A.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The LLM is used only for editing and formatting purposes in this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ Appendix C Implementation Details of the Token-level Decoding
This reformulation simplifies implementation by allowing a logit pre-processor to be applied before the Residual Aligner computes the $\mathrm{Softmax}$. This pre-processor retains only the tokens sampled from the Proposal Module, setting the logits for other tokens to $\mathrm{Inf}$, which is similar to the implementation of Nucleus Sampling. This adjustment enables the process to proceed through the standard $\mathrm{Softmax}$ and sampling procedure, allowing for effective token selection.
To mitigate performance degradation during the training of small Residual Aligner, $Q_{\theta}$, we only conduct secondary sampling when the distribution difference between the Proposal Module, $P_{\mathrm{M}}$, and the $Q_{\theta}$ is not significant. Specifically, we assess the difference using KL divergence $D_{KL}(P_{\mathrm{M}}\|Q_{\theta})$). If the KL divergence exceeds 0.1, indicating degradation of the Residual Aligner, we sample directly from the $P_{\mathrm{M}}$; otherwise, we apply the $Q_{\theta}$ for secondary sampling.